The volume of an ideal gas is 1 litre and it`s pressure is equal to 72 cm of mercury column . The volume of gas is made 900 cm3 by compressing it isothermally. The stess of the gas will be

Grade:12

1 Answers

Arun
6 years ago
Dear Abhishek
P1 * v1 = P2 * v2
72 * 1000 = P2 * 900
P2 = 80 cm( mercury)
Hence
Stress = \DeltaP = P2 - P1
 = 80 - 72 = 8cm (mercury)
Regards
Arun (askIITians forum expert)